De Nouvelles perspectives sur le Théorème de Morse-Sard

SEAN M. BATES et CARLOS G. MOREIRA

C.M.: IMPA, Estrada Dona Castorina 110, Jd. Botânico, 22460-320, Rio de Janeiro-RJ, Brasil. e-mail: gugu@impa.br

S.B.: Emerging Markets Trading, Deutschebank, 1 Great Winchester Street, London EC2N 2DB. e-mail: sean.bates@db.com

————————————————————————————

Résumé. Soient E un espace vectoriel de dimension finie, F un espace vectoriel normé, k ≥ 1 un entier, f: E → F une application de classe C^{k}, t > 0 un réel, et A ⊂ E un ensemble borné et de H^{t}-mesure finie, où H^{t} est la mesure de Hausdorff en dimension t. Si la différentielle de f en tout point de A est de rang ≤ r, nous montrons que l’image f(A) est de H^{d}-mesure nulle, avec d = r + \frac{t − r}{k} · Lorsque de plus D^{k}f satisfait une condition de Hölder d’exposant

α ∈ (0, 1], f(A) est de Hd̃-mesure finie, avec d̃= r + \frac{t − r}{k + α} ·

Lorsque E est un espace de Banach, les résultats précédents sont encore vrais si k ≤ 2. Par contre, nous présentons un polynôme P: ℓ^{2} → ℝ dont l’ensemble critique est de H^{3}-mesure finie et a pour image [0, 1].

Some new perspectives on Sard’s Theorem

Abstract. Let E be a vector space of finite dimension, F a normed vector space, k ≥ 1 an integer, f: E → F a C^{k} function, t > 0 a real number and A ⊂ E a bounded subset with finite H^{t}-measure, where H^{t} is the Hausdorff measure of dimension t. If the differential of f in every point of A has rank ≤ r, we show that the image f(A) has zero H^{d}-measure, where d = r + \frac{t − r}{k} · Moreover, if D^{k}f satisfies a Hölder condition of exponent α ∈ (0, 1], then f(A)

has finite Hd̃-measure, where d̃= r + \frac{t − r}{k + α} ·

If E is a Banach space, then the preceding results are still true if k ≤ 2. On the other side, we present a polynomial function P: ℓ^{2} → ℝ for which the critical set has finite H^{3}-measure, and whose image is [0, 1].

———————————————————————————— Abridged English Version

The classical Morse-Sard theorem states that the critical value set of a C^{k} mapping f: ℝ^{n} → ℝ^{m} has zero lebesgue measure in ℝ^{m}, provided that k ≥ max{n − m + 1, 1}. We shall give here the main ideas of the proof of the following generalizations of it:

Let t > 0 be a real number and X a metric space. We recall that a subset A ⊂ X is t-finite if H^{t}(A) < ∞, where H^{t} is the Hausdorff measure of dimension t, that A ⊂ X is t-null if H^{t}(A) = 0, and that A ⊂ X is t-rectifiable if t is integer and A is contained (mod. a t-null set) in a countable union of Lipschitz images of open sets of ℝ^{t}.

Let E be a vector space of finite dimension, F a normed vector space, k ≥ 1 an integer, α ∈ [0, 1] and f: E → F a C^{k+α} function (i.e., a k-times differentiable function such that D^{k}f satisfies a Hölder condition of exponent α).

Theorem 1: Under the above conditions, if t > 0 is a real number, r < t is an integer, A ⊂ E is a t-finite set such that rank(Df(x)) ≤ r, ∀ x ∈ A and d = r + (t − r)/(k + α), then:

(i) If α = 0 (i.e., f ∈ C^{k}), then f(A) is d-null.

(ii) If ||f|_{A}||_{C^{k+α}} < ∞ then there is a constant C = C(||f||_{C^{k+α}} , t, r, k, α) such that H^{d}(f(A)) ≤ C H^{t}(A).

(iii) If t is integer and A is t-rectifiable, then f(A) is d-null.

Remark: If α = 1 then statements (ii) and (iii) of the theorem remain true even if we just suppose that the k-th derivative of f satisfies a local Zygmund condition.

The proof of Theorem 1 uses induction on the dimension of the domain E of the function f. For differentiable mappings defined on infinite-dimensional spaces, the results of Theorem 1 are true if k + α ≤ 3. These results are obtained by substituting the inductive argument by a simpler argument based on the mean-value theorem, that works if k ≤ 2.

Theorem 2: Under the conditions of Theorem 1, if E is a Banach space then the same conclusions are true provided k ≤ 2, α ∈ [0, 1].

The proof of Theorem 1 is based on auxiliary results whose proof uses techniques similar to the techniques of [4]. For instance, in order to prove Theorem 1 (i), use the following

Lemma 1.i: Let η > 0 be a fixed constant, E a finite dimensional vector space, F a normed vector space and A ⊂ E a set of points of rank ≤ r of a differentiable function f: E → F such that ||f|_{A}||_{C^{1}} < ∞.

If k ≥ 1 is an integer and f ∈ C^{k} then there is a decomposition {A_{i}}_{i∈ℕ} of A and, for each i ∈ ℕ, a decomposition E = E′_{i} × E′′_{i}, with dim E′_{i} ≤ r such that if S ⊂ E,

diam(f(S ∩ A_{i})) ≤ (||f|_{A_{i}}||_{C^{1}} + η) diam(π_{E′_{i}}S) + η(diam S)^{k}.

We will use this lemma to prove Theorem 1.(i): We can take a piece A_{i} of the decomposition of A given by the lemma, and we may suppose that A_{i} ⊂ ⋃^{∞} ; i=1 C_{i}, where C_{i} is a cube of side ε_{i}

and ∑^{∞} ; i=1 ε_{i}^{t} ≤ 1. We take δ_{i} = ε_{i}^{k} and we divide C_{i} as the union of (ε_{i}/ηδ_{i})^{p} blocks of side ηδ_{i} in the direction of E′_{i}, where p = dim E′_{i} ≤ r. According to the lemma, if B is one of these blocks, diam(f(B ∩ A_{i})) ≤ (||f|_{A}||_{C^{1}} + η)ηδ_{i} + ηε_{i}^{k} ≤ Kηε_{i}^{k}, where K is a constant. Thus, we have

∑_{B⊂C_{i}} diam(f(B ∩ A_{i}))^{d} ≤ (ε_{i}^{1−k}/η)^{p}(Kηε_{i}^{k})^{j} ≤ K^{d}η^{d−p} ε_{i}^{kd+p(1−k)} = O(ε_{i}^{k}),

since p ≤ r, d > r and kd+r(1−k) = t. This shows that H^{d}(f(A_{i})) = 0, and so H^{d}(f(A)) = 0.

The results of the Theorems 1 and 2 are quite sharp. We can adapt the examples of [4] to show that we cannot relax the hypothesis of Theorem 1. Theorem 2 implies that for any Banach space E, a C^{2+ lips.} function f: E → ℝ maps any t-dimensional critical set to a set of dimension ≤ t/3. The following example shows that this result cannot be improved by improving the differentiability hypothesis:

Example: Let f: ℓ^{2} → ℝ given by

f( ∑_{n=1}^{∞} x_{n}e_{n}) = ∑_{n=1}^{∞}(3.2^{−nα} x_{n}^{2} − 2x_{n}^{3}), α ∈ [\frac{1}{3} , +∞).

Then f ∈ C^{∞} (moreover, f^{(4)} ≡ 0, so f is “polynomial”) and its critical set is C =

∏^{∞} ; n=1 {0, 2^{−nα}} = { ∑^{∞} ; n=1 x_{n}e_{n}, x_{n} ∈ {0, 2^{−nα}}}.

We have H^{1/α}(C) < +∞, and f(C) = { ∑^{∞} ; n=1 δ_{n} 2^{−3mα}, δ_{n} ∈ {0, 1}}, which is a Cantor set of

finite and positive measure of dimension 1/3α if α > \frac{1}{3} and is the interval [0, 1] if α = \frac{1}{3} · This shows that the conclusions of Theorem 1 are false if k ≥ 3.

—————————————–

Selon le théorème classique de Morse-Sard, l’ensemble des valeurs critiques d’une fonction f: ℝ^{n} ⟶^{C^{k}} ℝ^{m} est de mesure nulle en ℝ^{m} si k ≥ max{n − m + 1, 1}. La base de ce résultat est le lemme de décomposition suivant, de A.P. Morse:

Lemme de Criticalité de Morse: Soit A ⊂ ℝ^{n} l’ensemble critique d’une fonction f: ℝ^{n} ⟶^{C^{k}} ℝ. Il y a une décomposition dénombrable A = ⋃^{∞} ; i=1 A_{i} telle que pour chaque i ∈ ℕ et

chaque point d’accumulation x de l’ensemble A_{i} on a lim_{\begin{matrix}y→x \\ y∈A_{i} \end{matrix}} \frac{|f(y) − f(x)|}{|y − x|^{k}} = 0.

Ce lemme peut être utilisé pour démontrer que la dimension de Hausdorff de f(A) est majorée par n/k.

Nous allons présenter ici les idées principales de la démonstration des généralisations suivantes du théorème de Morse-Sard.

Soit t > 0 un réel et X un espace métrique. Un sous-ensemble A ⊂ X est dit t-fini si H^{t}(A) < ∞, où H^{t} est la mesure de Hausdorff en dimension t; A est t-nul si H^{t}(A) = 0 ; A est t-réctifiable (pour t entier) si A est (modulo un ensemble t-nul) contenu dans une union dénombrable d’images lipschitziennes d’ouverts de ℝ^{t}.

Soient E un espace vectoriel de dimension finie, F un espace vectoriel normé, k ≥ 1 un nombre entier, α ∈ [0, 1] et f: E → F une fonction de classe C^{k+α} (i.e., une fonction de classe C^{k} telle que D^{k}f satisfait une condition de Hölder d’exposant α).

Théorème 1: Sous les conditions précédentes, si t > 0 est un réel, r < t un entier, A ⊂ E est un ensemble t-fini tel que rang(Df(x)) ≤ r, ∀ x ∈ A et d = r + (t − r)/(k + α), alors:

(i) Si α = 0 (i.e., f ∈ C^{k}) alors H^{d}(f(A)) = 0.

(ii) Si ||f|_{A}||_{C^{k+α}} < ∞ alors il y a une constante C = C(||f||_{C^{k+α}} , t, r, k, α) telle que H^{d}(f(A)) ≤ CH^{t}(A).

(iii) Si t ∈ ℤ_{+} et A est t-rectifiable alors H^{d}(f(A)) = 0.

Remarque: Les conclusions (ii) et (iii) du Théorème 1 sont vraies dans le cas α = 1 même si nous supposons seulement que D^{k}f satisfait une condition de Zygmund locale.

Théorème 2: Lorsque E est un espace de Banach les conclusions du Théorème 1 son encore valables pour k ≤ 2 , α ∈ [0, 1].

La démonstration du Théorème 1 est basée sur le lemme suivant, qui est démontré en utilisant des techniques similaires aux techniques de [4].

Lemme 1: Soient η > 0, E un espace euclidien, F un espace vectoriel normé et A ⊂ E un ensemble de points de rang r d’une fonction différentiable f: E → F.

1) Si k ∈ ℤ_{+}, k ≥ 2 et f ∈ C^{k} alors il existe une décomposition {A_{i}} de A et, pour chaque i, une décomposition E = E′_{i} × E′′_{i} avec dim E′_{i} ≤ r telle que si S ⊂ E,

diam(f(S ∩ A_{i})) ≤ (∥f|_{A_{i}}∥_{C^{1}} + η)diam(π_{E′_{i}}S) + η(diam S)^{k}

2) Si k ∈ ℤ_{+}, α ∈ [0, 1] et k + α > 2 alors il existe une décomposition {A_{i}} de A telle que pour chaque i il y a une constante M_{i} > 0 et une décomposition E = E′_{i} × E′′_{i} avec dim E′_{i} ≤ r telle que si S ⊂ E,

diam(f(S ∩ A_{i})) ≤ (∥f|_{A_{i}}∥_{C^{1}} + η)diam(π_{E′_{i}}S) + M_{i}(diam S)^{k+α}

3) Si A est t-rectifiable pour t ∈ ℤ_{+}, alors il y a une décomposition {A_{i}} de A telle que H^{t}(A_{0}) = 0, ∑_{i}(diam A_{i})^{t} ≤ H^{t}(A) + 1 et pour chaque i ≠ 0 il y a une constante M_{i} > 0 et une décomposition E = E′_{i} × E′′_{i} avec dim E′_{i} ≤ r telle que si S ⊂ E,

diam(f(S ∩ A_{i})) ≤ (∥f|_{A_{i}}∥_{C^{1}} + η)diam(π_{E′_{i}}S) + η(diam A_{i})^{k+α}

Pour la démonstration du Théorème 2, ce lemme sera remplacé par le lemme suivante, qui peut être démontré en utilisant le théorème des accroissements finis:

Lemme 2: Soit E un espace de Banach. Si k ∈ {1, 2} et α ∈ [0, 1] alors les conclusions du Lemme 1 sont vraies.

On va démontrer le Théorème 1, (i). Pour cela, on va utiliser la partie 1) du Lemme 1. Les autres parties du Théorème 1 et le Théorème 2 suivent les mêmes idées, en utilisant les autres parties du Lemme 1 et le Lemme 2.

On peut prendre une pièce A_{j} de la décomposition de A donnée par le Lemme 1, (1), et on peut supposer que A_{j} ⊂ ⋃_{i=1}^{∞} C_{i}, où C_{i} est un cube de coté ε_{i} et ∑_{i=1}^{∞} ε_{i}^{t} ≤ 1. On prend δ_{i} = ε_{i}^{k} et on divise C_{i} dans la direction de E′_{j} en (ε_{i}/ηδ_{i})^{p} pavés de coté ηδ_{i} dans la direction de E′_{j}, où p = dim E′ ≤ r. Selon le lemme, si B est un de ces pavés, diam(f(B ∩ A_{j})) ≤ (||f||_{1} + η)ηδ_{i} + ηε_{i}^{k} ≤ Kηε_{i}^{k}, où K est une constante, et on a ∑_{B⊂C_{i}} diam(f(B ∩ A_{j}))^{d} ≤

(ε_{i}^{1−k}/η)^{p} · (Kηε_{i}^{k})^{d} ≤ K^{d}η^{d−p}ε_{i}^{kd+p(1−k)} = o(ε_{i}^{t}), parce que p ≤ r, d > r et kd + r(1 − k) = t. Cela démontre que H^{d}(f(A_{j})) = 0, d’où H^{d}(f(A)) = 0.

Les résultats des Théorèmes 1 et 2 sont les meilleurs possibles, comme on peut démontrer par des exemples, qui sont des variations de ceux de [4] pour le Théorème 1 et le suivant pour le Théorème 2.

Le Théorème 2 implique que pour un espace de Banach E et une fonction f: E → ℝ de classe C^{2+ lips.}, l’image d’un ensemble de dimension t est de dimension ≤ t/3. L’exemple suivant montre qu’on ne peut pas améliorer ce résultat en changeant les hypothèses de différentiablilité.

Exemple: Soit F: ℓ^{2} → ℝ donnée par

F ( ∑_{n=1}^{∞} x_{n}e_{n}) = ∑_{n=1}^{∞}(3 · 2^{−nα}x_{n}^{2} − 2x_{n}^{3}), α ∈ [\frac{1}{3}, +∞).

Alors F est de classe C^{∞} (nous avons même F^{(4)} ≡ 0, donc F est “polynomiale”), et

l’ensemble critique de F est C = ∏_{n=1}^{∞}{0, 2^{−nα}} = { ∑_{n=1}^{∞} x_{n}e_{n}, x_{n} ∈ {0, 2^{−nα}}}.

Nous avons H^{1/α}(C) < +∞, et F (C) = { ∑_{n=1}^{∞} δ_{n}2^{−3mα}, δ_{n} ∈ {0, 1}}, qui est un ensemble de Cantor de mesure finie et positive en dimension 1/3α si α > \frac{1}{3} et est égal a [0, 1] si α = \frac{1}{3} . Cela montre que les conclusions du Théorème 1 sont fausses pour k ≥ 3.

Références bibliographiques

[1] Bates, S.M., Toward a precise smoothness hypothesis in Sard’s theorem. Proc. Amer. Math. Soc. 117 (1993) 279–283.

[2] Bates, S.M. and Moreira, C.G., A generalization of Sard’s theorem, preprint, 2000.

[3] H. Federer, Geometric measure theore, Grundlehren Math. Wiss., vol. 153, Springer-Verlag, New York, (1969).

[4] Moreira, C., Hausdorff measures and the Morse-Sard theorem, preprint, submited to Publicacions Mathematiques.

[5] Morse, A., The Behavior of a function on its critical set, Ann. Math. 40 (1939) 62–70.

[6] Norton, A., The Zygmund Morse-Sard theorem, J. Geom. Analysis 4 (1994) 403–424.

[7] Sard, A., Images of critical sets, Ann. Math. 68 (1958) 247–259.

[8] Sard, A., Hausdorff measure of critical images on Banach manifolds, Amer. J. Math. 87 (1965), 158–174.